Let p be the population proportion for the following condition. Find the point estimates for p and q. Of 887 children​ surveyed, 115 plan to join the armed forces in the future. The point estimate for​ p, Modifying Above p with caret​, is __________​(Round to three decimal places as​ needed.)
Answer #1

here as parameter of interest is population proportion of children that join the armed forces in the future

point estimate for​ p =115/887=0.130

point estimate for​ q=1-0.130=0.870